# Scaling Laws Meet Model Architecture: Toward Inference-Efficient LLMs

### v1

@@ 1 Introduction
Finally, we validate this framework by fitting the proposed scaling law on models between 80M and 297M parameters, and evaluating its predictions when scaling up to 3B-parameter models. Our results demonstrate that, under identical training setups, the derived optimal 3B-parameter architecture achieves $42\%$ higher inference throughput than the LLaMA-3.2-3B architecture, while maintaining better accuracy.
@@ 3.1 Model Architecture Variations
Figure 2: Inference throughput vs (left) hidden size $d=d_{\text{model}}$ and (right) mlp-to-attention ratio $r=r_{\text{mlp}/\text{attn}}$ on the 8B model. Under a fixed parameter budget $N_{\text{non-embed}}$, larger hidden sizes and higher mlp-to-attention ratios improve inference throughput for varying batch sizes.
@@ 3.2 Inference Efficiency
Inspired by the success and widespread adoption of open-weight dense models such as Qwen3 (Yang et al., 2025), LLaMA-3.2 (Dubey et al., 2024), and the Gemma-2 (Team et al., 2024b) family, we construct architectural variants by modifying the configurations of the LLaMA-3.2 and Qwen3 dense models (Figure 11-13 in Appendix E). In addition to hidden size and the mlp-to-attention ratio, we find that group-query attention has a critical impact on inference efficiency, even though it only modestly reduces the number of attention parameters (by shrinking the key and value matrices). To disentangle these effects, we conduct controlled ablations of hidden size, MLP-to-attention ratio, and GQA under the following setups:
@@
  • GQA: fix $N_{\text{non-embed}}$, $d_{\text{model}}$ and $r_{\text{mlp}/\text{attn}}$, vary $n_{\text{head}}$ and number of key-value heads (Appendix E).
Figure 2 shows the ablation of varying hidden sizes $d_{\text{model}}$ and mlp-to-attention $r_{\text{mlp}/\text{attn}}$ on the LLaMA-3.1-8B model variants. We observe that larger hidden size (or fewer attention heads) and higher mlp-to-attention ratios improve inference throughput. Similar trends are observed in the LLaMA-3.2-1B and 3B model variants (Appendix E). These gains arise in part because larger $d_{\text{model}}$ and higher $r_{\text{mlp}/\text{attn}}$ reduce the total FLOPs, as detailed in the inference FLOPs analysis (Appendix H). In addition, these architectural choices shrink the KV cache, lowering I/O cost during inference and further improving throughput Adnan et al. (2024). Figure 10 in Appendix E presents the GQA ablation, confirming prior observations Ainslie et al. (2023) that increasing GQA consistently improves inference throughput. A comparable set of ablation experiments on Qwen3 models, also reported in Appendix E, further corroborates these findings.
@@ 3.3 A Conditional Scaling Law
Figure 3: Loss vs. hidden size: (Left) 80M model variants; (Center) 145M model variants; (Right) 297M model variants. Across model sizes, the relationship between training loss and $d_{\text{model}}/\sqrt{N}$ exhibits a consistent U-shaped curve when architectural factors such as GQA and the MLP-to-attention ratio are held fixed. The legend denotes the MLP-to-attention ratio $r=r_{\text{mlp}/\text{attn}}$ for each model.
Figure 4: Loss vs. MLP-to-attention ratio: (Left) 80M model variants; (Center) 145M model variants; (Right) 297M model variants. Across model sizes, the relationship between training loss and $r_{\text{mlp}/\text{attn}}$ exhibits a consistent U-shaped curve when architectural factors such as GQA and hidden size are held fixed. The legend denotes the hidden size $d=d_{\text{model}}$ for each model.
@@
As shown in Figures 3 and 4, both hidden size and the MLP-to-attention ratio exhibit U-shaped relationships with training loss. To capture these trends, we fit the function $c_{0}+c_{1}\log x+c_{2}/x$ separately for $x=r_{\text{mlp/attn}}$ and $d_{\text{model}}/\sqrt{N_{\text{non-embed}}}$. This formulation effectively models the U-shaped behavior while ensuring sublinear growth as $x$ increases. However, incorporating $r_{\text{mlp}/\text{attn}}$, $d_{\text{model}}$, $N$, and $D$ into a unified, architecture-aware scaling law remains challenging. Since fitting a single all-purpose scaling law $L(d/\sqrt{N},r,N,D)$ is unrealistic across all possible configurations, we instead propose a two-step conditional approach:
@@
We focus on two simple calibration schemes:
@@
Here, $a_{i}$ and $b_{i}$ are learnable parameters that are shared across all $N,D$. Unlike the unified formulation, the conditional scaling law assumes that the effects of $r_{\text{mlp}/\text{attn}}$ and $d_{\text{model}}$ on loss are separable. We further ablate joint, non-separable formulations in Appendix G, where we find that they yield inferior predictive performance.
@@ 3.4 Searching for Inference-Efficient Accurate Models
As shown in Figure 10 (Appendix E), GQA has a substantial impact on inference efficiency; However, unlike hidden size and the mlp-to-attention ratio, GQA does not exhibit a consistent relationship with loss (Figure 14) and is highly variable, making it challenging to identify settings that achieve both accuracy and efficiency. Fortunately, the search space for GQA is relatively small once $N_{\text{non-embed}}$, $d_{\text{model}}$, and $r_{\text{mlp}/\text{attn}}$ are fixed, since GQA must be a prime factor of the number of attention heads $n_{\text{head}}$. In practice, we perform a local GQA search by enumerating feasible values and applying early stopping once performance falls below that of the GQA=4=4 baseline. Algorithm 1 summarizes our overall framework for identifying inference-efficient and accurate architectures.
@@ 4 Experiment Setup
We evaluate the inference efficiency using the vLLM framework Kwon et al. (2023). By default, inputs consist of 4096 tokens and outputs of 1024 tokens. We report the averaged inference throughput (tokens/second) from 5 repeated runs. Unless otherwise specified, all experiments are conducted on NVIDIA Ampere A100 GPUs (40GB).
@@ 5 Experiment Results
The mlp-to-attention ratio $r_{\text{mlp}/\text{attn}}$ of open-weights models typically fall between $0.5$ and $5$, for example, the mlp-to-attention ratio for LLaMA-3.2-1B, LLaMA-3.2-3B, and Qwen3-8B are 4.81, 1.5, and 4.67, respectively. In Figure 5, we fit the conditional scaling law using only model architectures with $r_{\text{mlp/attn}}\in[0.5,5]$. We ablate this choice by training model architectures with outlier $r_{\text{mlp}/\text{attn}}$ below $0.5$ and above $5$ (such as $0.1,12.6$) in Appendix C. In Figure 15 (left) and Figure 15 (center) in Appendix G, we show on Task 3 a comparison of fitting the conditional scaling law without and with these outliers (with a clear Spearman correlation score degradation), which suggests to exclude extreme outliers for better predicted performances.
@@
In Figure 15 (right), We ablate an alternative formulation of the scaling laws with additive calibration, as discussed in §3.3. The results on Task 3 show that multiplicative and additive calibrations achieve similar MSE and Spearman correlations, underscoring the robustness of our two-step reference plus calibration framework.
Table 1: Large-Scale Model Results: We evaluate the scaling laws and framework at the 1B and 3B scales by training Panda-1B, Surefire-1B, and Panda-3B, and compare them with LLaMA-3.2-1B and LLaMA-3.2-3B, respectively. The Avg. column reports the mean accuracy across the nine downstream tasks. Panda-1B and Panda-3B are trained using the optimal architectural configurations predicted by our scaling laws, whereas Surefire-1B and Surefire-3B satisfy the loss constraint in Eq. (4) and achieve Pareto optimality.
@@
Figure 6: Results for 1B and 3B models: (left) Panda-1B closely follows the scaling law predictions for minimizing training loss. (center) Inference throughput comparison between LLaMA-3.2-1B and Surefire-1B, showing that Surefire-1B consistently achieves higher efficiency across batch sizes. (right) Inference throughput comparison between LLaMA-3.2-3B and Surefire-3B, demonstrating that Surefire-3B consistently delivers higher efficiency across all batch sizes.
@@ 5.1 Optimal Model Architecture
With all components in place, we apply the search framework for inference-efficient and accurate models (Alg. 1). For the $N_{\text{non-embed}}=1$B and 33B setting trained on 100100B tokens, we set the target loss $L_{t}$ to match the training loss achieved by the LLaMA-3.2-1B and LLaMA-3.2-3B architectures, respectively. Although inference efficiency $I_{N}(P)$ could, in principle, be expressed analytically, it depends heavily on hardware and inference configurations. Therefore, rather than solving for $I_{N}(P)$ directly, we search over feasible configurations $P_{i}$} that satisfy the loss constraint and select Pareto-optimal points, which we denote as Surefire-1B and Surefire-3B. Surefire-1B and Surefire-3B outperform LLaMA-3.2-1B and LLaMA-3.2-3B on downstream tasks (Table 1) and deliver up to 42% higher inference throughput (Figure 6, center and right). Detailed downstream task accuracies are provided in Appendix I.
Table 2: 3B Model Ablation Study: We evaluate the robustness of scaling laws at 3B scale by training Panda-3B and Panda-3B, and compare them with LLaMA-3.2-3B. The Avg. column reports the mean accuracy across the nine downstream tasks. Panda-3B represents the optimal architectural configuration predicted by the conditional scaling laws fitted using the 80M, 145M, and 297M model data, whereas Panda-3B corresponds to the optimal configuration predicted from scaling laws fitted using the 1B model data.
@@
Figure 7: Effect of the Fitting Dataset on Predictive Performance vs (left) Fit on 80, 145, 297M, 1B, evaluate on 3B; (right) Fit on 1B, evaluate on 3B. Orange dots denote fitting data points, and purple crosses indicate the test data points. We compare scaling-law predicted loss with actual pretraining loss of architectures and we observe that fitting the scaling laws with only 1B model data yields lower MSE and higher Spearman correlation for the 3B model loss prediction.
Ablation of fitting data. While we adopt a progressive strategy for selecting fitting data across tasks (§4), results from small models (e.g., 80M) may not reliably predict behavior at larger scales such as 3B. To examine this, we perform an ablation study by fitting the conditional scaling law for the 3B model using only results from the 1B variant. As shown in Figure 7, fitting with only the 1B data achieves a lower MSE and a higher Spearman correlation when predicting the 3B loss. This indicates that the coefficients of the conditional scaling law may shift as model size increases.
We therefore refit the law with multiplicative calibration using only the 1B variants, yielding the coefficients
$a_{0}=2.319,a_{1}=0.238,a_{2}=0.0176,b_{0}=0.5104,b_{1}=0.0051,\text{and}~b_{2}=0.0062.$
This produces an alternative optimal configuration for the 3B model, with $d_{\text{model}}/\sqrt{N}=0.074$ and $r=1.229$. We train a 3B model (Panda-3B) under this configuration on 100B tokens and compare it with both LLaMA-3.2-3B and Panda-3B (fitted from 80M, 145M, 297M, and 1B data). As shown in Table 2, Panda-3B achieves a lower training loss and comparable downstream accuracy to Panda-3B, with detailed results given in Appendix I. These findings suggest that when scaling up, it is often sufficient, and sometimes preferable, to fit the law using models within a closer size range to the target, such as about one third of its scale.

### v2

@@ Abstract
Figure 1: Scaling sweep results. (left) Inference throughput (tokens/s) and (right) Scaling-law-predicted training loss contours over hidden size $d_{\text{model}}$ and mlp-to-attention ratio. Our conditional scaling law enables concurrent gains in throughput and reductions in predicted training loss under a fixed parameter budget. Dotted points indicate the architectures used to fit the scaling law.
@@ 1 Introduction
Finally, we validate this framework by fitting the proposed scaling law on models between 80M and 297M parameters, and evaluating its predictions when scaling up to pretrain 3B-parameter models. Our results demonstrate that, under identical training setups, the derived optimal 3B-parameter architecture achieves up to $42\%$ higher inference throughput than the LLaMA-3.2-3B architecture, while maintaining better accuracy.
@@ 3.1 Model Architecture Variations
Figure 3: Inference throughput. (left) hidden size $d=d_{\text{model}}$ and (right) mlp-to-attention ratio $r=r_{\text{mlp}/\text{attn}}$ on the 8B model. Under a fixed parameter budget $N_{\text{non-embed}}$, larger hidden sizes and higher mlp-to-attention ratios improve inference throughput for varying batch sizes.
@@ 3.2 Inference Efficiency
Inspired by the success and widespread adoption of open-weight dense models such as Qwen3 (Yang et al., 2025), LLaMA-3.2 (Dubey et al., 2024), and the Gemma-2 (Team et al., 2024b) family, we construct architectural variants by modifying the configurations of the LLaMA-3.2 and Qwen3 dense models (Figure 12-14 in Appendix F). In addition to hidden size and the mlp-to-attention ratio, we find that group-query attention has a critical impact on inference efficiency, even though it only modestly reduces the number of attention parameters (by shrinking the key and value matrices). To disentangle these effects, we conduct controlled ablations of hidden size, MLP-to-attention ratio, and GQA under the following setups:
@@
  • GQA: fix $N_{\text{non-embed}}$, $d_{\text{model}}$ and $r_{\text{mlp}/\text{attn}}$, vary $n_{\text{head}}$ and number of key-value heads (Appendix F).
Figure 3 shows the ablation of varying hidden sizes $d_{\text{model}}$ and mlp-to-attention $r_{\text{mlp}/\text{attn}}$ on the LLaMA-3.1-8B model variants. We observe that larger hidden size (or fewer attention heads) and higher mlp-to-attention ratios improve inference throughput. Similar trends are observed in the LLaMA-3.2-1B and 3B model variants (Appendix F). These gains arise in part because larger $d_{\text{model}}$ and higher $r_{\text{mlp}/\text{attn}}$ reduce the total FLOPs, as detailed in the inference FLOPs analysis (Appendix K). In addition, these architectural choices shrink the KV cache, lowering I/O cost during inference and further improving throughput Adnan et al. (2024). Figure 11 in Appendix F presents the GQA ablation, confirming prior observations Ainslie et al. (2023) that increasing GQA consistently improves inference throughput. A comparable set of ablation experiments on Qwen3 models, also reported in Appendix F, further corroborates these findings.
@@ 3.3 A Conditional Scaling Law
Figure 4: Loss vs. hidden size. (left) 80M model variants; (center) 145M model variants; (right) 297M model variants. Across model sizes, the relationship between training loss and $d_{\text{model}}/\sqrt{N}$ exhibits a consistent U-shaped curve when architectural factors such as GQA and the MLP-to-attention ratio are held fixed. The legend denotes the MLP-to-attention ratio $r=r_{\text{mlp}/\text{attn}}$ for each model.
Figure 5: Loss vs. MLP-to-attention ratio. (left) 80M model variants; (center) 145M model variants; (right) 297M model variants. Across model sizes, the relationship between training loss and $r_{\text{mlp}/\text{attn}}$ exhibits a consistent U-shaped curve when architectural factors such as GQA and hidden size are held fixed. The legend denotes the hidden size $d=d_{\text{model}}$ for each model.
@@
As shown in Figures 4 and 5, both hidden size and the MLP-to-attention ratio exhibit U-shaped relationships with training loss. To capture these trends, we fit the function $c_{0}+c_{1}\log x+c_{2}/x$ separately for $x=r_{\text{mlp/attn}}$ and $d_{\text{model}}/\sqrt{N_{\text{non-embed}}}$. This formulation effectively models the U-shaped behavior while ensuring sublinear growth as $x$ increases. However, incorporating $r_{\text{mlp}/\text{attn}}$, $d_{\text{model}}$, $N$, and $D$ into a unified, architecture-aware scaling law remains challenging. Since fitting a single all-purpose scaling law $L(d/\sqrt{N},r,N,D)$ is unrealistic across all possible configurations, we instead propose a two-step conditional approach:
@@
We focus on two simple and transparent calibration schemes:
@@
Here, $a_{i}$ and $b_{i}$ are learnable parameters that are shared across all $N,D$. Note that both functional forms assume the effects of $r_{\text{mlp}/\text{attn}}$ and $d_{\text{model}}$ on loss are separable.
@@ 3.4 Searching for Inference-Efficient Accurate Models
As shown in Figure 11 (Appendix F), GQA has a substantial impact on inference efficiency; However, unlike hidden size and the mlp-to-attention ratio, GQA does not exhibit a consistent continuous relationship with loss (Figure 24, Appendix I) and is highly variable, making it challenging to identify settings that achieve both accuracy and efficiency. Fortunately, the search space for GQA is relatively small once $N_{\text{non-embed}}$, $d_{\text{model}}$, and $r_{\text{mlp}/\text{attn}}$ are fixed, since GQA must be a prime factor of the number of attention heads $n_{\text{head}}$. In practice, we perform a local GQA search by enumerating feasible values and applying early stopping once performance falls below that of the GQA=4=4 baseline. Algorithm 1 summarizes our overall framework for identifying inference-efficient and accurate architectures.
@@ 4 Experiment Setup
We evaluate the inference efficiency using the vLLM framework Kwon et al. (2023). By default, inputs consist of 4096 tokens and outputs of 1024 tokens. We report the averaged inference throughput (tokens/second) from 5 repeated runs. Unless otherwise specified, all experiments are conducted on NVIDIA Ampere A100 GPUs (40GB) with vLLM.
@@ 5 Experiment Results
The mlp-to-attention ratio $r_{\text{mlp}/\text{attn}}$ of open-weights models typically fall between $0.5$ and $5$, for example, the mlp-to-attention ratio for LLaMA-3.2-1B, LLaMA-3.2-3B, and Qwen3-8B are 4.81, 1.5, and 4.67, respectively. In Figure 6, we fit the conditional scaling law using only model architectures with $r_{\text{mlp/attn}}\in[0.5,5]$. We ablate this choice by training model architectures with outlier $r_{\text{mlp}/\text{attn}}$ below $0.5$ and above $5$ (such as $0.1,12.6$) in Appendix D. In Figure 25 (left) and Figure 25 (center) in Appendix J, we show on Task 3 a comparison of fitting the conditional scaling law without and with these outliers (with a clear Spearman correlation score degradation), which suggests to exclude extreme outliers for better predicted performances.
@@
In Figure 25 (right), we ablate an alternative formulation of the scaling laws with additive calibration, as discussed in §3.3. The results on Task 3 show that multiplicative and additive calibrations achieve similar MSE and Spearman correlations. Note that, unlike the conventional unified formulation, both calibrations assume that the effects of $r_{\text{mlp}/\text{attn}}$ and $d_{\text{model}}$ on loss are separable. We further ablate more complex joint, non-separable formulations in Appendix J and find that they do not provide superior predictive performance. The two-step reference-and-calibration framework appears robust enough that simple calibrations perform well.
Table 1: Large-Scale Model Results. We evaluate the scaling laws at 1B and 3B scales by training Panda-1B, Surefire-1B, and Panda-3B, and compare them with LLaMA-3.2-1B and LLaMA-3.2-3B, respectively. The Avg. column reports the mean accuracy across the nine downstream tasks. Panda-1B and 3B are trained using the optimal architectural configurations predicted by our scaling laws, whereas Surefire-1B and 3B satisfy the loss constraint in Eq. (4) and achieve Pareto optimality.
@@
Figure 7: Results for 1B and 3B models. (left) Panda-1B closely follows the scaling law predictions for minimizing training loss. (center & right) Inference throughput comparison between LLaMA-3.2 and Surefire models, where Surefire is consistently efficient across all batch sizes.
@@ 5.1 Optimal Model Architecture
With all components in place, we apply the search framework for inference-efficient and accurate models (Alg. 1). For the $N_{\text{non-embed}}=1$B and 33B setting trained on 100100B tokens, we set the target loss $L_{t}$ to match the training loss achieved by the LLaMA-3.2-1B and LLaMA-3.2-3B architectures, respectively.
Ablation of inference efficiency. Although inference efficiency $I_{N}(P)$ could, in principle, be expressed analytically, it depends heavily on hardware and inference configurations. Therefore, rather than solving for $I_{N}(P)$ directly, we search over feasible configurations $P_{i}$} that satisfy the loss constraint on A100 with vLLM and select Pareto-optimal points, which we denote as Surefire-1B and Surefire-3B. Surefire-1B and Surefire-3B outperform LLaMA-3.2-1B and LLaMA-3.2-3B on downstream tasks (Table 1 with details in Appendix L) and deliver up to 42% higher inference throughput (Figure 7, center and right). We also ablate inference efficiency using both vLLM and SGLang Zheng et al. (2023) on A100 and NVIDIA H200 GPUs (Appendix FG). The results remain consistent with our vLLM–A100 evaluation: Surefire-1B and 3B outperform LLaMA-3.2-1B and 3B across all settings, achieving up to 47% higher throughput with SGLang on H200. This demonstrates that the efficiency gains transfer across serving stacks and hardware platforms. Detailed throughput statistics are provided in Table 6.
Table 2: 3B Model Ablations. We assess the robustness of fitting-data strategy at 3B scale by training Panda-3B (using 80M, 145M, and 297M data) and Panda-3B (using only on 1B data), and compare both with LLaMA-3.2-3B. Avg. denotes mean accuracy across nine downstream tasks.
@@
Figure 8: Effect of the Fitting Data Strategy on Predictive Performance. (left) Fit on 80M, 145M, 297M, 1B, evaluate on 3B; (right) Fit on 1B, evaluate on 3B. Orange dots denote fitting data, and purple crosses indicate the test data. We compare scaling-law predicted loss with actual pretraining loss of architectures and we observe that fitting the scaling laws with only 1B model data yields lower MSE and higher Spearman correlation for the 3B model loss prediction.
Ablation of fitting data strategy. While we adopt a progressive strategy for selecting fitting data across tasks (§4), results from small models (e.g., 80M) may not reliably predict behaviors at larger scales such as 3B. To assess this, we fit the conditional scaling law for the 3B model using only the 1B variants. As shown in Figure 8, fitting with 1B data yields lower MSE and higher Spearman correlation when predicting 3B behavior, suggesting that the law’s coefficients shift with model size. We therefore refit the law with multiplicative calibration using only the 1B variants, yielding the coefficients $a_{0}=2.319$, $a_{1}=0.238$, $a_{2}=0.0176$, $b_{0}=0.5104$, $b_{1}=0.0051$, and $b_{2}=0.0062$.
This produces an alternative optimal configuration for the 3B model, with $d_{\text{model}}/\sqrt{N}=0.074$ and $r=1.229$. We train a 3B model (Panda-3B) under this configuration on 100B tokens and compare it with both LLaMA-3.2-3B and Panda-3B (fitted from 80M, 145M, 297M, and 1B data). As shown in Table 2, Panda-3B achieves a lower training loss and comparable downstream accuracy to Panda-3B, with detailed results given in Appendix L. These findings suggest that when scaling up, it is often sufficient, and sometimes preferable, to fit the law using models within a closer size range to the target, such as about one third of its scale.
@@ 6 Related Work
Scaling laws are powerful tools to predict the performance of large language models. Existing scaling laws Hoffmann et al. (2022); Muennighoff et al. (2023); Sardana et al. (2023); Kumar et al. (2024); Gadre et al. (2024); Ruan et al. (2024) characterize how model performance varies with model size, dataset size, data quality, and compute budget. With the rise of Mixture-of-Experts (MoE) Shazeer et al. (2017); Guo et al. (2025), a powerful architecture for large language models, recent studies Krajewski et al. (2024); Abnar et al. (2025) extend scaling laws to account for the number of experts, expert granularity, active parameters, and sparsity. Due to space constraints, we defer additional related work to Appendix B.
7 Limitations and Future Work
While our team has made notable progress, several open challenges remain that offer promising directions for future research. First, due to limitations in resources and time, our evaluation does not extend to 7B models. Second, our analysis is restricted to dense models, and it remains unclear whether the results extend to Mixture of Experts (MoE) architectures Shazeer et al. (2017). While we report inference efficiency measurements for MoE models under varying architectural choices in Appendix M, we have not yet established scaling laws for MoE architectures. Finally, our analysis is limited to pre-training, and it remains unclear how the results would change under post-training.
8 Conclusion
This work explores the trade-off between model accuracy and inference cost under a fixed training budget. We begin by demonstrating how architectural choices influence both inference throughput and model accuracy. Building on this, we extend Chinchilla scaling laws to incorporate architectural factors and propose a two-step conditional framework for optimal architecture search: (i) train small models to fit the conditional scaling law (Eq. 3), and (ii) solve Eq. 4 for the predicted optimal architecture, followed by a local search over GQA to maximize inference efficiency. Using the fitted scaling laws and our framework, we trained models up to 3B parameters, achieving up to 42% higher inference throughput and 2.1% accuracy gains across nine downstream tasks. In Table 7 and Table 8 of Appendix H, we compare design choices across existing open-source models at the 1B and 3B scales, further underscoring the need for our inference-efficient accurate model designs.
Reproducibility Statement
All experiments in this work were conducted using publicly available frameworks. Section 4 provides details of our training, inference, and evaluation setups. In particular, we used Megatron-LM (Shoeybi et al., 2019) for model training, vLLM (Kwon et al., 2023) and SGLang (Zheng et al., 2023) for efficient inference, and lm-eval-harness (Gao et al., 2024a) for standardized evaluations.
Acknowledgements
Song Bian and Shivaram Venkataraman acknowledge the support of the NSF Diamond project OAC-2311767 (Democratizing Large Neural Network Model Training for Science).
Appendix A LLM Usage
Appendix B Additional Related Work